OpenStudy (anonymous):

simplify. (1+(14/a))/(1-14/a))

OpenStudy (anonymous):

\[(1+14/a)/(1-14/a)\]

OpenStudy (anonymous):

i got an answer of 1 :\

OpenStudy (anonymous):

Is it \[\frac{a+14}{a-14}\]

OpenStudy (anonymous):

can you show me how? did you distribute an a or something?

OpenStudy (anonymous):

\[(\frac{a}{a}+\frac{14}{a})\div(\frac{a}{a}-\frac{14}{a})\]

OpenStudy (anonymous):

\[(\frac{14+a}{a})\div(\frac{a-14}{a})\]

OpenStudy (anonymous):

ahh. then you flipped the second part right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

great, thanks!